# TwinVLA: Data-Efficient Bimanual Manipulation with Twin Single-Arm Vision-Language-Action Models

### v1

@@ Abstract
Vision-language-action models (VLAs) trained on large-scale robotic datasets have demonstrated strong performance on manipulation tasks, including bimanual tasks. However, because most public datasets focus on single-arm demonstrations, adapting VLAs for bimanual tasks typically requires substantial additional bimanual data and fine-tuning. To address this challenge, we introduce TwinVLA, a modular framework that composes two copies of a pretrained single-arm VLA into a coordinated bimanual VLA. Unlike monolithic cross-embodiment models trained on mixtures of single-arm and bimanual data, TwinVLA improves both data efficiency and performance by composing pretrained single-arm policies. Across diverse bimanual tasks in real-world and simulation settings, TwinVLA outperforms a comparably-sized monolithic RDT-1B model without requiring any bimanual pretraining. Furthermore, it narrows the gap to state-of-the-art model, $\pi_{0}$ which rely on extensive proprietary bimanual data and compute cost. These results establish our modular composition approach as a data-efficient and scalable path toward high-performance bimanual manipulation, leveraging public single-arm data.
@@ 1 Introduction
Thanks to publicly available large-scale robotic datasets, vision-language-action models (VLAs) have shown impressive performance in single-arm robotic manipulation, generalizing across diverse tasks, objects, and environments (rt22023arxiv; oxe2024; kim2024openvla; black2024pi_0). However, extending these successes to bimanual manipulation remains challenging, as public bimanual datasets are scarce, and existing approaches often rely on large, proprietary datasets that require thousands of hours of data collection and curation (black2024pi_0), limiting reproducibility and progress.
Can we build strong bimanual VLAs without collecting or fine-tuning on large bimanual datasets by leveraging existing single-arm data? Recent cross-embodiment learning work typically trains monolithic models on multi-robot datasets (oxe2024) by employing embodiment-specific action decoders (octo_2023; doshi2024scalingcrossembodiedlearningpolicy; nvidia2025gr00tn1openfoundation) or shared, zero-padded action spaces (liu2024rdt; black2024pi_0). Although promising, differences in observation and action spaces introduce heterogeneity, forcing a single model to handle disparate action spaces, and monolithic training underutilizes the modular structure inherent to bimanual tasks.
A modular perspective on bimanual manipulation is supported by neuroscience: human bimanual manipulation is the coordination of arm-specific motor primitives rather than a single monolithic controller. Dedicated neural circuits, such as the Supplementary Motor Area (SMA) and the corpus callosum, orchestrate and synchronize the two arms (Sadato9667; Swinnen2002). Similar principles have proven effective in vision-language modeling, where interaction between modality-specific backbones improves its efficiency and effectiveness (liang2024mixtureoftransformerssparsescalablearchitecture).
@@
Inspired by these insights, we propose TwinVLA, a modular architecture that operationalizes this coordination-centric view. Instead of training from scratch, TwinVLA leverages a pretrained single-arm VLA. Specifically, we first design a lightweight, compact single-arm VLA, which we call SingleVLA (Appendix˜A). We pretrain SingleVLA for single-arm manipulation on the OXE dataset (oxe2024). We then duplicate this SingleVLA and integrate the two “twin” instances through a lightweight coordination method. This design is highly data-efficient: it eliminates the need for a bimanual pretraining dataset and achieves strong performance with only a small amount of bimanual demonstrations for fine-tuning.
To integrate two SingleVLAs into a bimanual policy, TwinVLA utilizes a joint attention (liang2024mixtureoftransformerssparsescalablearchitecture) across the twin models, as illustrated in Figure˜1. This allows the twin SingleVLAs to exchange information and coordinate their actions, while preserving their pretrained capabilities. This approach is made feasible without significant overhead, as we duplicate only the VLM backbone and utilize a Mixture-of-Experts (MoE) to efficiently manage shared inputs. In contrast to monolithic cross-embodiment models (liu2024rdt; octo_2023; doshi2024scalingcrossembodiedlearningpolicy), our approach yields better performance and data efficiency, significantly reducing the need for large-scale bimanual data collection and compute.
@@
Figure 2: (a) Data efficiency. While RDT-1B and $\pi_{0}$ use 1M + single-arm data with sizable bimanual data, TwinVLA uses $\sim$0.5M single-arm data and only 50 target bimanual data. (b) Compute efficiency. RDT-1B and $\pi_{0}$ require high compute (exceeding 1,000 H100 GPU-days), whereas TwinVLA achieves higher or comparable performance at substantially lower compute (25 H100 GPU-days).
@@ 2 Related Work
Vision-Language-Action models (VLAs) represent a promising approach to robotic control. They adapt large pretrained Vision-Language Models (VLMs) (liu2023visualinstructiontuning; karamcheti2024prismaticvlmsinvestigatingdesign; chen2024internvlscalingvisionfoundation) to robotic control (oxe2024; khazatsky2024droid) by fine-tuning on action-labeled data. This process enables them to translate language instructions and visual context into grounded actions. Early VLA models explore various strategies for connecting VLMs with action generation (ahn2022icanisay). RT-2 (rt22023arxiv) tokenizes actions as part of the model’s output vocabulary, whereas RoboFlamingo (li2024visionlanguagefoundationmodelseffective) uses a separate continuous action head. Recent work, such as $\pi_{0}$ (black2024pi_0) and CogAct (li2024cogactfoundationalvisionlanguageactionmodel), further integrates a denoising policy with a VLM to improve action execution accuracy.
@@
Bimanual manipulation policies are essential for enabling robots to perform complex tasks that require coordinated two-handed control, such as folding laundry (bersch2011bimanual; avigal2022speedfolding), assembling parts (stavridis2018bimanual), or wiping the plate black2025pi; chi2024universalmanipulationinterfaceinthewild. Learning effective bimanual policies is challenging due to high-dimensional, tightly coupled action spaces and the scarcity of high-quality bimanual demonstrations (lee2020learning; xie2020deepimitationlearningbimanual). Consequently, specialist methods, such as Diffusion policy (chi2023diffusion) and ACT (zhao2023learning), trained only on target-task demonstrations, struggle on precise, long-horizon bimanual tasks (black2024pi_0).
To mitigate data scarcity, prior work has largely pursued a straightforward approach: training a single, unified model through large-scale bimanual data collection and computationally intensive pretraining. RDT-1B (liu2024rdt) is fine-tuned on over 66K bimanual episodes after extensive pretraining on the large-scale OXE dataset (oxe2024), reportedly requiring a month on $48$ H100 GPUs. Similarly, $\pi_{0}$ (black2024pi_0) relies on a $10,000$-hour proprietary dataset with substantial computational cost. Moreover, because their data are not publicly accessible, reproducibility and broader adoption are limited.
@@
In contrast to such monolithic cross-embodiment policies and compute-heavy bimanual pretraining, our approach adopts a modular, coordination-centric design. We first train a SingleVLA on large-scale public single-arm data, duplicate the pretrained SingleVLA and couple them via joint attention, and then fine-tune it on bimanual tasks—allowing each stage to benefit from the most suitable data (see Figure˜2). This composition-based approach avoids bimanual pretraining, requires only a small amount of bimanual fine-tuning, better preserving the capabilities of single-arm policies, and improves data and compute efficiency.
@@ 3 Preliminaries
This paper aims to develop a data-efficient framework for learning bimanual manipulation policies by building upon pretrained single-arm Vision-Language-Action (SingleVLA) models. This section formalizes the single-arm and bimanual settings, and describes the conditional flow-matching objective for training action heads of VLAs.
@@ 3.1 Formulating the Bimanual VLA Policy
Our goal is to extend a pretrained SingleVLA $\pi_{\text{single}}$ into a bimanual policy $\pi_{\text{twin}}$ applicable to target bimanual tasks. A VLA $\pi(A_{t}\mid o_{t})$ predicts an action chunk $A_{t}=(a_{t},a_{t+1},\dots,a_{t+T-1})$ of length $T$ from an observation $o_{t}$. For single-arm manipulation, the observation $o^{\text{{single}}}_{t}=\left((l,I_{\text{ego}})_{t},(I_{\text{wrist}},d)_{t}\right)$ includes a language prompt $l$, a ego-centric image $I_{\text{ego}}$ (shared input), and an arm-specific wrist image $I_{\text{wrist}}$ with proprioception $d$ (arm-specific input). We train $\pi_{\text{single}}(A_{t}\mid o^{\text{single}}_{t})$ to predict the action chunk for one arm. For bimanual manipulation, the observation aggregates both right ($R$) and left ($L$) arm-specific input, $o^{\text{twin}}_{t}=\left((l,I_{\text{ego}})_{t},(I_{\text{wrist}}^{R},d^{R})_{t},(I_{\text{wrist}}^{L},d^{L})_{t}\right)$, and the policy $\pi_{\text{twin}}(A^{R}_{t},A^{L}_{t}\mid o^{\text{twin}}_{t})$ outputs a joint action chunk for right and left arms.
@@ 3.2 Training VLAs with Conditional Flow Matching
We aim for the VLA model to predict continuous robot actions from observations. Each observation $o_{t}$ is tokenized: language via a language tokenizer, images via a vision encoder, and proprioception via an MLP encoder. We append a learnable readout token $r_{t}$ to the observation token sequence, which signals the model to produce actions. These tokens are fed into the VLM backbone, and the embedding $h_{t}$ corresponding to the readout token $r_{t}$ is obtained from the final hidden state.
To enable continuous action prediction from VLM outputs, we attach an action head using conditional flow matching for both $\pi_{\text{single}}$ and $\pi_{\text{twin}}$. The action head $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$ is trained with the following flow-matching loss function:
@@
where $h_{t}$ is the VLM output hidden state, $d_{t}$ is proprioception. The objective trains the action head to predict the reference flow from a noised action chunk $A_{t}^{\tau}$ to the target action chunk $A_{t}$, where $\tau\in[0,1]$ denotes the flow matching timesteps. We adopt a simple linear Gaussian path $q(A_{t}^{\tau}\mid A_{t})=N(\tau A_{t},(1-\tau)I)$, which has demonstrated robust performance across domains.
Specifically, we first sample a noisy action $A_{t}^{\tau}=\tau A_{t}+(1-\tau)\epsilon$, where the noise $\epsilon\sim N(0,I)$ and the timestep $p(\tau)=Beta(\frac{0.999-\tau}{0.999};1.5,1)$, following $\pi_{0}$ (black2024pi_0). The action head $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$ and the VLM are jointly trained end-to-end to predict the reference flow $u(A_{t}^{\tau}\mid A_{t})=\epsilon-A_{t}$.
During inference, we sample actions using the forward Euler integration method. Starting from $A_{0}\sim N(0,I)$, we iteratively update the action using the learned flow $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$:
@@
where we set the sampling step $n=10$ and use $\delta=\frac{1}{n}$. The entire model, including the vision encoder and VLM backbone, is trained in an end-to-end manner.
@@ 4 TwinVLA
TwinVLA is a modular architecture that transforms a pretrained SingleVLA into a coordinated bimanual policy through three core principles. First, we selectively duplicate modules from a pretrained SingleVLA to form a bimanual policy (Section˜4.1). We then introduce a joint attention mechanism to enable effective cross-arm coordination between the duplicated VLMs (Section˜4.2). Finally, we integrate a Mixture-of-Experts (MoE) layer for shared observations to enhance computational efficiency without sacrificing performance (Section˜4.3).

### v2

@@ Abstract
Vision-language-action models (VLAs) trained on large-scale robotic datasets have demonstrated strong performance on manipulation tasks, including bimanual tasks. However, because most public datasets focus on single-arm demonstrations, adapting VLAs for bimanual tasks typically requires substantial additional bimanual data and fine-tuning. To address this challenge, we introduce TwinVLA, a modular framework that composes two copies of a pretrained single-arm VLA into a coordinated bimanual VLA. Unlike monolithic cross-embodiment models trained on mixtures of single-arm and bimanual data, TwinVLA improves both data efficiency and performance by composing pretrained single-arm policies. Across diverse bimanual tasks in real-world and simulation settings, TwinVLA outperforms a comparably-sized monolithic RDT-1B model without requiring any bimanual pretraining. Furthermore, it narrows the gap to state-of-the-art model $\pi_{0}$, which relies on extensive proprietary bimanual data and compute cost. These results establish our modular composition approach as a data-efficient and scalable path toward high-performance bimanual manipulation, leveraging public single-arm data.
@@ 1 Introduction
Thanks to publicly available large-scale robotic datasets, vision-language-action models (VLAs) have shown impressive performance in single-arm robotic manipulation, effectively adapting to downstream tasks and generalizing across diverse tasks, objects, and environments (Zitkovich et al., 2023; Open X-Embodiment Collaboration et al., 2024; Kim et al., 2024; Black et al., 2024). However, extending these successes to bimanual manipulation remains challenging, as public bimanual datasets are scarce, and existing approaches often rely on large, proprietary datasets that require thousands of hours of data collection and curation (Black et al., 2024), limiting reproducibility and progress.
Can we build strong bimanual VLAs without collecting or fine-tuning on large bimanual datasets by leveraging existing single-arm data? In this work, we propose a highly data-efficient adaptation paradigm for bimanual control that eliminates the need for prohibitive bimanual pretraining. By effectively repurposing a single-arm VLA, we demonstrate that complex bimanual skills can be mastered using only minimal target-domain demonstrations, establishing a practical and reproducible pathway to bimanual manipulation.
To effectively realize this transfer of single-arm priors to bimanual control, the choice of underlying architecture is critical. Recent cross-embodiment learning work typically trains monolithic models on multi-robot datasets (Open X-Embodiment Collaboration et al., 2024) employing embodiment-specific action decoders (Octo Model Team et al., 2024; NVIDIA et al., 2025) or shared, zero-padded action spaces (Liu et al., 2024; Black et al., 2024). Although promising, differences in observation and action spaces introduce heterogeneity, forcing a single model to handle disparate action spaces, and monolithic training underutilizes the modular structure inherent to bimanual tasks.
A modular perspective on bimanual manipulation is supported by neuroscience: human bimanual manipulation is the coordination of arm-specific motor primitives rather than a single monolithic controller. Dedicated neural circuits, such as the Supplementary Motor Area (SMA) and the corpus callosum, orchestrate and synchronize the two arms (Sadato et al., 1997; Swinnen, 2002). Similar principles have proven effective in vision-language modeling, where interaction between modality-specific backbones improves its efficiency and effectiveness (Liang et al., 2024).
@@
Inspired by these insights, we propose TwinVLA, a modular architecture that operationalizes this coordination-centric view. Instead of training from scratch, TwinVLA leverages a pretrained single-arm VLA. Specifically, we first design a lightweight, compact single-arm VLA, which we call SingleVLA (Appendix˜A). We pre-train a 0.8B-size SingleVLA for single-arm manipulation on the OXE dataset (Open X-Embodiment Collaboration et al., 2024). We then duplicate this SingleVLA and integrate the two “twin” instances through a lightweight coordination method. This design is highly data-efficient: it eliminates the need for a bimanual pretraining dataset and achieves strong performance with only a small amount of bimanual demonstrations for fine-tuning.
To integrate two SingleVLAs into a bimanual policy, TwinVLA utilizes a joint attention (Liang et al., 2024) across the twin models, as illustrated in Figure˜1. This allows the twin SingleVLAs to exchange information and coordinate their actions, while preserving their pretrained capabilities. This approach is made feasible without significant overhead, as we duplicate only the VLM backbone and utilize a Mixture-of-Experts (MoE) to efficiently manage shared inputs. In contrast to monolithic cross-embodiment models (Liu et al., 2024; Octo Model Team et al., 2024; Doshi et al., 2024), our approach yields better performance and data efficiency, significantly reducing the need for large-scale bimanual data collection and compute.
@@
Figure 2: (a) Data efficiency. TwinVLA requires only $\sim 800$h of single-arm and $50$ episodes of target bimanual data, significantly less than RDT-1B ($\sim 2,400$h) and $\pi_{0}$ ($\sim 10,900$h) in total. (b) Compute efficiency. RDT-1B and $\pi_{0}$ require high compute (exceeding $1,000$ H100 GPU-days), whereas TwinVLA achieves higher or comparable performance with only $25$ H100 GPU-days.
@@ 2 Related Work
Bimanual manipulation policies are essential to enable robots to perform complex tasks that require coordinated two-handed control, such as folding laundry (Bersch et al., 2011; Avigal et al., 2022), assembling parts (Stavridis and Doulgeri, 2018), or wiping the plate (Black et al., 2025; Chi et al., 2024b). Learning effective bimanual policies is challenging due to high-dimensional, tightly coupled action spaces and the scarcity of high-quality bimanual demonstrations (Lee et al., 2020; Xie et al., 2020). Consequently, specialist methods, such as Diffusion policy (Chi et al., 2024a) and ACT (Zhao et al., 2023), trained only on target-task demonstrations, struggle on precise, long-horizon tasks.
Recent works have explored various architectures for bimanual control to explicitly model the inter-dependencies between arms (Lee et al., 2024; Kobayashi and Buamanee, 2025), or focus on high-level language planning via VLM (Gbagbe et al., 2024). Anybimanual (Lu et al., 2025) introduces a high-level skill manager to coordinate primitives and visual aligner to mask 3D voxels for decoupled policies, benefiting from both high-level managing and architectural inductive bias. While promising, it is difficult to generalize these methods, as they are often limited to small-scale scenarios, handle only low-dexterity tasks, or backbone constraints (Grotz et al., 2024; Shridhar et al., 2022).
Alternatively, another line of research extends successful unimanual Vision-Language-Action (VLA) models (Liu et al., 2023b; Zitkovich et al., 2023; Li et al., 2024b) to bimanual tasks. This transition is challenging due to the scarcity of bimanual data, as public datasets are predominantly unimanual. To overcome this, prior work trains ‘monolithic’ models, requiring large-scale bimanual data collection and intensive pretraining. For example, RDT-1B (Liu et al., 2024) required massive pretraining and fine-tuning (reportedly a month on $48$ H100 GPUs), and $\pi_{0}$ (Black et al., 2024) relies on a $10,000$-hour proprietary dataset, both incurring high computational costs. Furthermore, the proprietary nature of these datasets limits reproducibility and broader adoption.
@@
In contrast to both monolithic, compute-heavy pretraining and specialized architectural designs, our approach adopts a modular, coordination-centric design. While Anybimanual (Lu et al., 2025) introduces novel inductive biases for coordination, these are often difficult to integrate into general-purpose VLA frameworks due to specific backbone constraints. Our method, however, is designed to leverage and scale the existing generalist VLAs. We first train a SingleVLA on large-scale public single-arm data, duplicate to couple them, and then fine-tune it on bimanual tasks—allowing each stage to benefit from the most suitable data (see Figure˜2). This composition-based approach avoids bimanual pretraining, requires only a small amount of bimanual fine-tuning, better preserves the strong capabilities of single-arm policies, and significantly improves data and compute efficiency.
@@ 3 Preliminaries
This paper aims to develop a data-efficient framework for learning bimanual manipulation policies by building upon pretrained single-arm Vision-Language-Action (SingleVLA) models. This section formalizes the single-arm and bimanual settings, briefly describes the VLA training objective, and introduces the core architectural concepts we leverage.
@@ 3.1 Formulating the Bimanual VLA Policy
Our goal is to extend a pretrained SingleVLA $\pi_{\text{single}}$ into a bimanual policy $\pi_{\text{twin}}$ applicable to target bimanual tasks. A VLA $\pi(A_{t}\mid o_{t})$ predicts an action chunk $A_{t}=(a_{t},a_{t+1},\dots,a_{t+T-1})$ of length $T$ from an observation $o_{t}$. For single-arm manipulation, the observation $o^{\text{{single}}}_{t}=\left((l,I_{\text{ego}})_{t},(I_{\text{wrist}},d)_{t}\right)$ includes a language prompt $l$, an ego-centric image $I_{\text{ego}}$ (shared input), and an arm-specific wrist image $I_{\text{wrist}}$ with proprioception $d$ (arm-specific input). We train $\pi_{\text{single}}(A_{t}\mid o^{\text{single}}_{t})$ to predict the action chunk for one arm. For bimanual manipulation, the observation aggregates both right ($R$) and left ($L$) arm-specific input, $o^{\text{twin}}_{t}=\left((l,I_{\text{ego}})_{t},(I_{\text{wrist}}^{R},d^{R})_{t},(I_{\text{wrist}}^{L},d^{L})_{t}\right)$, and the policy $\pi_{\text{twin}}(A^{R}_{t},A^{L}_{t}\mid o^{\text{twin}}_{t})$ outputs a joint action chunk for right and left arms.
@@ 3.2 Training VLAs with Conditional Flow Matching
We train our VLA models to predict continuous robot actions from observations. Each observation $o_{t}$ is tokenized and fed into the VLM backbone to produce an output embedding $h_{t}$ (from a learnable readout token $r_{t}$). To enable continuous action prediction from $h_{t}$, we attach an action head $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$ and train it using a conditional flow matching objective. The action head is trained with the following loss function:
@@
where $h_{t}$ is the VLM output embedding and $d_{t}$ is proprioception. This objective trains the action head $v_{\theta}$ to predict the reference flow $\mathbf{u}$ from a noised action chunk $A_{t}^{\tau}$ to the target action chunk $A_{t}$, conditioned on the VLM output and proprioception.
During inference, we sample actions using the forward Euler integration method. Starting from $A_{0}\sim N(0,I)$, we iteratively update the action using the learned flow $v_{\theta}$:
@@
where we set the sampling step $n=10$ and use $\delta=\frac{1}{n}$.
3.3 Mixture-Based Architectures
To adapt Transformers for multi-modal inputs, various mixture-based architectures have been explored and shown to be effective. These approaches range from combining entire, modality-specific backbones to ensembling or mixing individual layers within a single backbone. We briefly introduce two such paradigms that inform our design: a model-level Mixture-of-Transformers (MoT), which coordinates separate backbones, and a layer-level Mixture-of-Experts (MoE), which enables efficient, sparse computation.
The MoT architecture (Liang et al., 2024) enables efficient information sharing between separate, modality-specific backbones (e.g., text and image). It introduces joint attention, a shared self-attention layer performed over the union of multimodal inputs, allowing each modality to directly attend to the others. Meanwhile, modality-specific components such as feed-forward networks remain separate, making fusion lightweight yet effective.
MoE (Shazeer et al., 2017) scales model capacity efficiently by routing each input $x$ through a weighted combination of expert feed-forward networks using a gating function, yielding $\text{MoE}(x)=\sum_{i}w_{i}E_{i}(x)$, where $w_{i}$ denotes the routing weight.
@@ 4 TwinVLA
TwinVLA is a modular architecture that transforms a pretrained single-arm VLA into a coordinated bimanual policy. The overall computation flow of our architecture is described in Algorithm˜1 and Figure˜3. TwinVLA integrates single-arm policies through three core principles: (1) selective module duplication (Section˜4.1), (2) cross-arm fusion via joint attention (Section˜4.2), and (3) efficient shared representation via Mixture-of-Experts (Section˜4.3).
Algorithm 1 TwinVLA
$X^{m}_{0}$: encoded inputs from $o_{t}^{\text{twin}}$ (Section˜3.1) for each input $m\in\{\text{shared},\text{left},\text{right}\}$.
$\text{FFN}_{b}$: feed-forward network layer from each backbone $b\in\{\text{left},\text{right}\}$.
NN: Number of transformer layers
for $n=0$ to N1N-1 do $\triangleright$ Iterate every transformer layer
   // Prepare Q, K, V for each input $m$
  for each input $m\in\{\text{shared},\text{left},\text{right}\}$ do
   $Q^{m}_{n},K^{m}_{n},V^{m}_{n}\leftarrow{\color[rgb]{0.1796875,0.546875,0.33984375}\text{Norm}(\text{Proj}(X^{m}_{n}))}$ $\triangleright$ Input-specific projections, Algorithm˜3
  end for
   // Joint attention across inputs with attention re-weighting
  $\{A^{m}_{n}\}\leftarrow{\color[rgb]{0.078125,0.4375,0.79296875}\text{JointAttention}(\{Q^{m}_{n}\},\{K^{m}_{n}\},\{V^{m}_{n}\},M)}$ $\triangleright$ Algorithm˜2, with mask $\color[rgb]{0.078125,0.4375,0.79296875}M$ Figure˜3(a)
   // Residual & FFN / MoE
  for each input $m\in\{\text{shared},\text{left},\text{right}\}$ do
   $H^{m}_{n}\leftarrow X^{m}_{n}+{\color[rgb]{0.1796875,0.546875,0.33984375}\text{Norm}(\text{Proj}(A^{m}_{n}))}$ $\triangleright$ Input-specific output projection, Algorithm˜3
   $F^{m}_{n}\leftarrow{\color[rgb]{1,0.12109375,0.04296875}\text{MoE}(H^{m}_{n})}$ if $m=\text{shared}$ else $\text{FFN}_{m}(H^{m}_{n})$ $\triangleright$ MoE for shared input, Equation˜3
   $X^{m}_{n+1}\leftarrow H^{m}_{n}+{\color[rgb]{0.1796875,0.546875,0.33984375}\text{Norm}(F^{m}_{n})}$ $\triangleright$ Residual connection with norm, Algorithm˜3
  end for
end for
return $\{X^{m}_{N}\}$ $\triangleright$ Return outputs, this will be used for action decoding
4.1 Single-Arm Policy Duplication
We first pre-train a VLA on a single-arm dataset, which we refer to as SingleVLA. Note that existing pre-trained models can also be used for this purpose. To construct TwinVLA from SingleVLA, we initialize the twin policies for the left and right arms by copying the pretrained SingleVLA. However, instead of duplicating the full model, we share the vision encoder and DiT (Peebles and Xie, 2023) action head while fully replicating the VLM. Each arm has its own lightweight proprioception encoder. This design yields a compact 1.31.3B-parameter model, comparable to the 1.21.2B-parameter RDT-1B, without significantly increasing computational cost.
4.2 Joint Attention for Cross-arm Fusion
We integrate arm-specific inputs using a Joint Attention mechanism inspired by MoT (Liang et al., 2024). As illustrated in Figure˜3(b) and Algorithm˜1, this is achieved by sharing only the self-attention layers across the VLM backbones. Specifically, we concatenate the $\color[rgb]{0.078125,0.4375,0.79296875}Q,K,V$ from both backbones, perform self-attention, and subsequently split the outputs back to their respective streams, while other components such as projections use arm-specific networks from each arm’s VLM backbone. Unlike $\pi_{0}$ (Black et al., 2024), which links a VLM with an action head, we connect two VLMs directly. We elaborate joint attention mechanism in detail on Algorithm˜2.